OpenStudy (anonymous):

How would you evaluate using the squeeze theorem? lim x→0 , x2/sin^2*9x

OpenStudy (blockcolder):

\[0\le\frac{1}{\sin^2{9x}}\le1\\ 0\le\frac{x^2}{\sin^2{9x}}\le x^2\]